#2
between two quantities can be established. Procedure: Use the following sets of data and work with each one. The equation for a linear graph is y mx+b, where m is the slope and b is the y-intercept DATA SET 1: Fahrenheit-vs-Celsius Fahrenheit Celsius 32 68 104 140 176 0 20 40 60 80 1. Using Data Set 1 above, graph Fahrenheit (y)-vs- Celsius (x), using the scatterplot function in excel, or another graphing software. Make sure you label the axes. Fahrenheit should be on the y-axis, and Celsius should be on the x-axis. Make sure you label the axes on your graph. 2. Include the linear equation in the plot. This is the trendline in excel. If calculating by hand, make sure to indicate on the plot which two points are being used. They should be points on the line, not data points. Write down the equation of the line below, and use it to find the slope of the line, and the y- intercept. Equation of the line: Slope of the line: Y-intercept: 3. The equation of the line should be the relationship between Fahrenheit and Celsius: °F (9/5)°C +32, with the slope equal to 9/5, and the y-intercept equal to 32. Does your data show this?
Answer #1

200 T 1S0 160 140 120 S0 60 40 10 20 30 40 50 70 S0

b)

m,slope=(176-32)/(80-0)=1.8=9/5   

(y-32) / (x-0)=9 / 5

y=(9x/5) +32 (A-ans)

m,slope=(176-32)/(80-0)=1.8=9/5 (B- ans)

for y intercept put x=0

y=32 (C-Ans)

c) by putting all values of Fahernheit in y and celcius in x we proves that the data is valid according to the equation also the equation we formed using our data reveals the same thing.